OpenStudy (anonymous):

Simplify the expression: \[16a^4+24a^2b^2+9b^4\]

OpenStudy (anonymous):

I figured it out.

OpenStudy (anonymous):

\[=16a^4+12a^2b^2+12a^2b^2+9b^4\] \[=4a^2(4a^2+3b^2)+3b^2(4a^2+3b^2)\] \[=(4a^2+3b^2)(4a^2+3b^2)\] \[=(4a^2+3b^2)^2\]